Determine values of the constants A and B such that an = An + B is a solution of


recurrence relation an = 2an−1 + n + 5. Hence, find the solution of this recurrence


relation with a0 = 4.

1
Expert's answer
2021-12-07T09:08:20-0500

Let us determine values of the constants "A" and "B" such that "a_n = An + B" is a particular solution of the recurrence relation "a_n = 2a_{n\u22121} + n + 5."

It follows that

"An+B=2(A(n-1)+B)+n+5=2An-2A+2B+n+5"

"=(2A+1)n+(2B-2A+5),"

and thus

"A=2A+1" and "B=2B-2A+5."

Therefore, "A=-1" and "B=2A-5=-7."

We conclude that "a_n=-n-7" is a paticular solution of the recurrence relation.


Further, let us find the general solution of this recurrence relation with "a_0 = 4."

It follows that the characteristic equation "k-2=0" has the solution "k=2," and hence the general solution is of the form "a_n=C\cdot2^n-n-7." Since "4=a_0=C-7," we conclude that "C=11."

Consequently, the general solution of this recurrence relation "a_n = 2a_{n\u22121} + n + 5" with "a_0 = 4" is

"a_n=11\cdot2^n-n-7."